# Progressive Prompt Detailing for Improved Alignment in Text-to-Image Generative Models

### v2

@@ Abstract
Figure 1: SCoPE (ours) vs Stable Diffusion-2.1 [17] for long, detailed prompts. Note how SCoPE (right) captures details mentioned in the prompt better compared to Stable Diffusion-2.1. Some details clearly missing from Stable Diffusion-2.1 but captured in SCoPE generations are underlined and marked in blue.
@@ 1 Introduction
To address this limitation,  Zhang et al. [21] extend the context length of the text encoder, allowing for better representation of longer prompts. While effective, this approach requires retraining on large-scale datasets, making it computationally expensive. Another line of work focuses on mitigating misalignment in the latent space [23] by conditioning on individual concepts sequentially at different stages of the denoising process. However, this method primarily focuses on concept misalignment and entanglement in short prompts and does not focus on addressing the challenges with longer, detailed prompts.
In this work, we propose SCoPE which stands for Scheduled interpolation of Coarse-to-fine Prompt Embeddings. SCoPE is a training-free approach that improves alignment between the provided (long) prompt and the generated image in diffusion models. Our key idea is to dynamically break down the input prompt into a series of sub-prompts starting from a coarse-grained description that captures the global scene layout to more fine-grained details. We draw inspiration from the findings in Park et al. [11] that diffusion denoising is a progressive coarse-to-fine generation process, where initial timesteps establish low-frequency, global structures, while later steps introduce high-frequency, fine-grained details. Specifically, while all prior methods rely on a single static embedding of the entire input prompt, SCoPE interpolates between progressively detailed prompt embeddings throughout the denoising process, thus generating the global scene layout before gradually introducing finer-grained details. We extensively evaluate SCoPE against several open-sourced models and show that SCoPE improves prompt alignment for long complex prompts (obtained from the GenAI-bench dataset [6]) (see Fig. 1) and achieves a +8 improvement in VQA-based text-image alignment scores over Stable Diffusion [17, 12] baselines. Notably, SCoPE is both training-free and easily extensible, requires minimal computational overhead of only $+0.7$ seconds per inference on $1$ A6000 GPU.
@@ 2 Related work
1 Robots do yoga.
2 Robots perform yoga in a sunny garden.
3 Colorful robots do yoga in a sunlit garden with flowers, lit by golden hour light.
4 A circle of colorful robots practicing yoga in a sunny garden with flowers and grass, illuminated by golden hour light.
5 A vibrant circle of colorful robots gracefully practicing yoga poses in a lush, sunlit garden, surrounded by blossoming flowers and gentle grasses. The scene is illuminated by soft golden hour light, with intricate details showcasing each robot’s unique design, creating a serene and harmonious atmosphere. $\dagger$
Table 1: Progressively detailed sub-prompts derived from the GenAI-Bench prompt a circle of robots doing yoga in a garden.. $\dagger$ denotes the final prompt used to generate the baseline image (having $56$ tokens). Refer to Fig. 3 for generation results.
@@ 3.1 Sub-prompt generation and interpolation schedule
Sub-prompt generation. We first use GPT-4o-mini [10] to break down a given prompt into $n$ progressively detailed sub-prompts, each depicting the same scene with increasing level of detail (an example in Table 1). We then obtain the CLIP embeddings [13] of each sub-prompt such that $\mathbf{p}_{1}$ corresponds to the embedding of the coarsest prompt and $\mathbf{p}_{n}$ to the final fine-grained prompt.
@@ 3.2 Interpolation-based text conditioning
[…]
Adjusting the coarse-to-fine transition in SCoPE. As denoising progresses, the width of the Gaussian assigned to each sub-prompt gradually narrows due to the time-dependent decay of $\sigma_{t}$, reducing the influence of coarser sub-prompts at later timesteps. This results in a gradual shift in conditioning from coarse to fine details. The hyperparameter $\sigma_{0}$ further modulates this transition. Higher $\sigma_{0}$ values allows sub-prompts to maintain their influence longer, thereby yielding a more gradual transition. By contrast, lower values of $\sigma_{0}$ makes coarse sub-prompts to lose influence more quickly and resulting in a sharper transition to fine details. We note that for timesteps $t>q_{n}$, $I(\mathbf{p},t)=\mathbf{p}_{n}$, i.e., interpolation is no longer applied, and the model conditions solely on the final fine-grained prompt embedding. This ensures that the image generation process is fully guided by the most detailed prompt, focusing on refining fine-grained details during the later denoising steps, during the fidelity-improvement phase, as discussed in Liu et al. [8]. Thus, the hyperparameter $q_{n}$ controls when fine-grained details begin to influence the denoising process, determining how early these details start to guide image generation. Fig. 3 shows how $\sigma_{0}$ and $q_{n}$ impact image generation.
Figure 3: Effect of $\sigma_{0}$ and $q_{n}$ on image generation. The figure illustrates how the variations in the initial standard deviation ($\sigma_{0}$) and interpolation period ($q_{n}$) influence the generated images and their VQA Scores [7]. Smaller $q_{n}$ values (e.g., 4) preserve fine details (e.g., flowers, unique robot designs), while larger $q_{n}$ values (e.g., 20) emphasize broader structural elements, such as scene composition (e.g., soft golden hour light) and object interactions (e.g., robots doing yoga). Similarly, smaller $\sigma_{0}$ values lead to images retaining more fine-grained details. See Sec. 4 for more details. The sub-prompts used for this example are provided in Table 1.
@@ 3.3 Evaluation Setup
Dataset: We evaluate our approach using prompts derived from GenAI-Bench [6], which contains 1600 prompts with tags representing spatial relation, counting, negation, etc. These prompts have an average token length of $14.8$, and often lack fine-grained details. To ensure we test our model on longer, more detailed prompts, we applied the method from Deckers et al. [2] and adopt prompt enhancement, and to target a length of around $50$ words, to align within 75 tokens of the CLIP Text encoder. This step results in increasing the average token length to $69.3$. Specifically, we use GPT-4o-mini [10] to generate a more detailed version of each prompt. Subsequently, we then use GPT-4o-mini to decompose the enhanced prompt to return four variations, each capturing the same scene at a different level of detail. The prompts used to generate both the enhanced and simplified versions are provided in the appendix. We use the final fine-grained prompt to generate the baseline image generations and to obtain the evaluation scores described next.
@@ 6.2 Simplification Prompt
Figure 5: More examples to compare SCoPE generated images and the images generated from the baseline Stable Diffusion-2.1. The hyperparameters ($q_{n}$, $\sigma_{0}$) used are mentioned on top of each image generated by SCoPE.

### v3

@@ 1 Introduction
Figure 1: SCoPE (ours) vs SDXL [12] for long, detailed prompts. Note how SCoPE (right) captures details mentioned in the prompt better compared to SDXL.
To address this limitation,  Zhang et al. [21] extend the context length of the text encoder, allowing for better representation of longer prompts. While effective, this approach requires retraining on large-scale datasets, making it computationally expensive. Another line of work focuses on mitigating misalignment in the latent space [23] by conditioning on individual concepts sequentially at different stages of the denoising process. However, this method primarily focuses on concept misalignment and entanglement in short prompts and does not focus on addressing the challenges with longer, detailed prompts. In this work, we propose SCoPE which stands for Scheduled interpolation of Coarse-to-fine Prompt Embeddings. SCoPE is a training-free approach that improves alignment between the provided (long) prompt and the generated image in diffusion models. Our key idea is to dynamically break down the input prompt into a series of sub-prompts starting from a coarse-grained description that captures the global scene layout to more fine-grained details. We draw inspiration from the findings in Park et al. [11] that diffusion denoising is a progressive coarse-to-fine generation process, where initial timesteps establish low-frequency, global structures, while later steps introduce high-frequency, fine-grained details. Specifically, while all prior methods rely on a single static embedding of the entire input prompt, SCoPE interpolates between progressively detailed prompt embeddings throughout the denoising process, thus generating the global scene layout before gradually introducing finer-grained details. We extensively evaluate SCoPE against several open-sourced models and show that SCoPE improves prompt alignment for long complex prompts (obtained from the GenAI-bench dataset [6]) (see Fig. 1) and achieves a +8 improvement in VQA-based text-image alignment scores over Stable Diffusion [17, 12] baselines. Notably, SCoPE is both training-free and easily extensible, requires minimal computational overhead of only $+0.7$ seconds per inference on $1$ A6000 GPU.
@@ 2 Related work
1 A cat rides a skateboard on a city street.
2 A cat on a skateboard navigates a city street, balancing amidst people and cars.
3 An earless cat rides a skateboard in a busy street, casting shadows, maintaining balance against a city backdrop.
4 A sleek, earless cat skillfully rides a colorful skateboard on a bustling urban street, highlighted by sunlight shadows, showing agility and balance against a vibrant city backdrop.
5 A sleek, earless cat gracefully rides a colorful skateboard amidst a bustling urban street, effortlessly navigating between pedestrians and vehicles. Sunlight casts dramatic shadows, highlighting the feline’s agile, streamlined form as it maintains perfect balance, capturing a sense of motion and boldness against the vibrant city backdrop. $\dagger$
Table 1: Progressively detailed sub-prompts derived from the GenAI-Bench prompt A cat without visible ears is riding. $\dagger$ denotes the final prompt used to generate the baseline image. Refer to Fig. 3 for generation results.
@@ 3.1 Sub-prompt generation and interpolation schedule
Sub-prompt generation. We first use GPT-4o [10] to break down a given prompt into $n$ progressively detailed sub-prompts, each depicting the same scene with increasing level of detail (an example in Table 1). We then obtain the CLIP embeddings [13] of each sub-prompt such that $\mathbf{p}_{1}$ corresponds to the embedding of the coarsest prompt and $\mathbf{p}_{n}$ to the final fine-grained prompt.
@@ 3.2 Interpolation-based text conditioning
After defining the interpolation schedule, we use it to apply a Gaussian-based weighting mechanism at each denoising timestep $t\leq q_{n}$. Specifically, we define a Gaussian of standard deviation $\sigma$ centered at $q_{i}$. This aligns with our motivation where early timesteps benefit from broader, coarse guidance, while later timesteps favor sharper focus on fine-grained details. We define a weight to assign to each prompt embedding $\mathbf{p}_{i}$ at denoising timestep $t$ as $\alpha_{i,t}=\exp\left(-\frac{(t-q_{i})^{2}}{2\sigma^{2}}\right)$, where the denominator $\sigma$ controls the sharpness of the Gaussian function. Following the symmetric decay of the Gaussian, during denoising, at each timestep $t\leq q_{n}$, weights assigned to earlier sub-prompts gradually decrease, while those for later sub-prompts increase. The weights $\alpha_{i,t}$ are then normalized to obtain $\alpha_{i,t}^{\prime}=\frac{\alpha_{i,t}}{\sum_{j=1}^{n}\alpha_{j,t}}$. The final text embedding to condition at each timestep is computed as a weighted sum of sub-prompt embeddings, i.e., $I(\mathbf{p},t)=\lVert\mathbf{p}_{n}\rVert\sum_{i=1}^{n}\alpha_{i,t}^{\prime}{\mathbf{\hat{p}}}_{i}$, where $\hat{\mathbf{p}}_{i}=\frac{\mathbf{p}_{i}}{\lVert\mathbf{p}_{i}\rVert}$. The rescaling of magnitude by $\lVert\mathbf{p}_{n}\rVert$ is performed to ensure that the interpolated embedding lies on the hypersphere defined by the CLIP embedding space.
Coarse-to-fine transition in SCoPE. As denoising progresses, influence of coarser sub-prompts at later timesteps decreases. This results in a gradual shift in conditioning from coarse to fine details. The hyperparameter $\sigma$ further modulates this transition. Higher $\sigma$ values allows sub-prompts to maintain their influence longer, thereby yielding a more gradual transition. By contrast, lower values of $\sigma$ makes coarse sub-prompts to lose influence more quickly and resulting in a sharper transition to fine details. We note that for timesteps $t>q_{n}$, $I(\mathbf{p},t)=\mathbf{p}_{n}$, i.e., interpolation is no longer applied, and the model conditions solely on the final fine-grained prompt embedding. This ensures that the image generation process is fully guided by the most detailed prompt, focusing on refining fine-grained details during the later denoising steps, during the fidelity-improvement phase, as discussed in Liu et al. [8]. Thus, the hyperparameter $q_{n}$ controls when fine-grained details begin to influence the denoising process, determining how early these details start to guide image generation. Fig. 3 shows how $\sigma$ and $q_{n}$ impact image generation.
Figure 3: Effect of $\sigma$ and $q_{n}$ on image generation. The figure illustrates how the variations in standard deviation ($\sigma$) and interpolation period ($q_{n}$) influence the generated images and their VQAScores. Smaller $q_{n}$ values (e.g., 4) preserve fine details (e.g., vehicles, pedestrians), while larger $q_{n}$ values (e.g., 20) emphasize broader structural elements, such as scene composition (e.g., cat, skateboard) and object interactions (e.g., riding). Similarly, smaller $\sigma$ values lead to images retaining more fine-grained details. See Sec. 4 for more details. The sub-prompts used for this example are provided in Table 1
@@ 3.3 Evaluation Setup
Dataset: We evaluate our approach using prompts derived from GenAI-Bench [6], which contains 1600 prompts with tags representing spatial relation, counting, negation, etc. These prompts have an average token length of $14.8$, and often lack fine-grained details. To ensure we test our model on longer, more detailed prompts, we applied the method from Deckers et al. [2] and adopt prompt enhancement, and to target a length of around $50$ words, to align within 75 tokens of the CLIP Text encoder. This step results in increasing the average token length to $69.3$. Specifically, we use GPT-4o [10] to generate a more detailed version of each prompt. Subsequently, we then use GPT-4o to decompose the enhanced prompt to return four variations, each capturing the same scene at a different level of detail. The prompts used to generate both the enhanced and simplified versions are provided in the appendix. We use the final fine-grained prompt to generate the baseline image generations and to obtain the evaluation scores described next.
@@ 5 Conclusion
More Qualitative Results
Figure 5: More examples to compare SCoPE generated images and the images generated from the baseline Stable Diffusion-2.1.